OpenStudy (anonymous):

There are some Formulas related to Hyperbolic Functions and Inverse Hyperbolic Functions:

OpenStudy (anonymous):

\[\large \color{green}{\textbf{HYPERBOLIC FUNCTIONS:}}\] \[\large 1.\quad sinhx = \frac{e^x-e^{-x}}{2}\] \[\large 2. \quad coshx = \frac{e^x + e^{-x}}{2}\] \[\large 3. \quad tanhx = \frac{e^x - e^{-x}}{e^x + e^{-x}}\] \[\large 4. \quad cothx = \frac{e^x+ e^{-x}}{e^x - e^{-x}}\] \[\large 5. \quad sechx = \frac{2}{(e^x + e^{-x})}\] \[\large 6. \quad cosechx = \frac{2}{(e^x -e^{-x})}\] \[\large \color{green}{\textbf{RELATIONS:}}\] \[\large 1. \quad tanhx = \frac{sinhx}{coshx}\] \[\large 2. \quad cothx = \frac{coshx}{sinhx}\] \[\large 3. \quad sechx = \frac{1}{coshx}\] \[\large 4. \quad cosechx = \frac{1}{sinhx}\] \[\large 5. \quad \cosh^2x - \sinh^2x = 1\] \[\large 6. \quad 1 - \tanh^2x = sech^2x\] \[\large 7. \quad \coth^2x - 1 = cosech^2x\] \[\large \color{green}{\textbf{NEGATIVE ANGLE PROPERTIES:}}\] \[\large 1. \quad \sinh(-x) = -sinhx\] \[\large 2. \quad \cosh(-x) = coshx\] \[\large 3. \quad \tanh(-x) = -tanhx\] \[\large \color{green}{\textbf{DOUBLE AND TRIPLE ANGLES FORMULAS:}}\] \[\large 1. \quad \sinh2x = 2sinhx.coshx\] \[\large 2. \quad \cosh2x = \cosh^2x + \sinh^2x\] \[\large 3. \quad \tanh2x = \frac{2tanhx}{1 + \tanh^2 x}\] \[\large 4. \quad \sinh3x = 3sinhx + 4\sinh^3x\] \[\large 5. \quad \cosh3x = 4\sinh^3x -3coshx\] \[\large 6. \quad \tanh3x = \frac{3tanhx + \tanh^3x}{1 + 3\tanh^2x}\] \[\large \color{green}{\textbf{SUM AND DIFFERENCE OF ANGLES:}}\] \[\large 1. \quad \sinh(\alpha \pm \beta) = \sinh \alpha.\cosh \beta \pm \cosh \alpha.\sinh \beta\] \[\large 2. \quad \cosh(\alpha \pm \beta) = \cosh \alpha.\cosh \beta \pm \sinh \alpha.\sinh \beta\] \[\large 3. \quad \tanh(\alpha \pm \beta) = \frac{\tanh \alpha \pm \tanh \beta}{1 \pm \tanh \alpha.\tanh \beta}\] \[\large \color{green}{\textbf{SIMPLE VERSUS HYPERBOLIC:}}\] \[\large 1. \quad cosix = coshx\] \[\large 2. \quad sinix = i.sinhx\] \[\large 2. \quad tanix = i.tanhx\] \[\large \color{green}{\textbf{INVERSE HYPERBOLIC FUNCTIONS:}}\] \[\large 1. \quad \sinh^{-1}x = Log_e(x + \sqrt{x^2 + 1})\] \[\large 2. \quad \cosh^{-1}x = Log_e(x + \sqrt{x^2 - 1})\] \[\large 3. \quad \tanh^{-1}x = \frac{1}{2}Log_e (\frac{1+x}{1-x})\] \[\large 4. \quad \coth^{-1}x = \frac{1}{2}Log_e (\frac{x+1}{x-1})\] \[\large 5. \quad sech^{-1}x = Log_e (\frac{1\pm \sqrt{1-x^2}}{x})\] \[\large 6. \quad cosech^{-1}x = Log_e (\frac{1+ \sqrt{1+x^2}}{x}), \quad x > 0\] \[\large = Log_e (\frac{1- \sqrt{1+x^2}}{x}), \quad x < 0\] \[\large \color{green}{\textbf{RELATIONS IN INVERSE HYPERBOLIC:}}\] \[\large 1. \quad \sinh^{-1}x = \cosh^{-1} \sqrt{1 + x^2}\] \[\large 2. \quad \cosh^{-1}x = \sinh^{-1} \sqrt{x^2 - 1}\]

mathslover (mathslover):

Very useful thanks :)

OpenStudy (anonymous):

Welcome @mathslover

OpenStudy (lgbasallote):

\[\huge \color{purple}{\textbf{<tips hat>}}\]

OpenStudy (anonymous):

I have seen you every time you write it @lgbasallote What it means??

OpenStudy (lgbasallote):

http://en.wikipedia.org/wiki/Hat_tip

OpenStudy (anonymous):

Ha ha ha ha..

OpenStudy (lgbasallote):

it's an act i've been doing for 7 months now as a show of respect to people :p

OpenStudy (anonymous):

Oh I got it now.. I have seen it in the picture in the link.. Ha ha ha..

OpenStudy (anonymous):

What are \(\mathsf{\text{Hyperbolic Functions}}\)?

OpenStudy (anonymous):

Do you know about simple Trigonometric Functions?? Also called Circular Trigonometric Functions...

OpenStudy (apoorvk):

I appreciate this @waterineyes , but could please help me understand what hyperbolic functions are actually?

OpenStudy (anonymous):

These are just same as Normal or Circular Trigonometric Functions, But these are expressed in exponential terms.. You can they are analogues of trigonometric functions..

OpenStudy (apoorvk):

Thanks, just one small example? Or may be a link where that explains this a bit lucidly? Thanks a bunch!

OpenStudy (anonymous):

I have one pdf with me may be it will help you:

OpenStudy (apoorvk):

Thank you!!!

OpenStudy (anonymous):

Welcome dear.. If you have any doubt then let me know that..

OpenStudy (unklerhaukus):

@waterineyes can you please explain these a bit: \[ \color{green}{\textbf{SIMPLE VERSUS HYPERBOLIC:}}\]\[1. \quad \cos ix = \cosh x\]\[2. \quad \sin ix = i\cdot\sinh x\]\[ 2. \quad \tan ix = i\cdot\tanh x\]

OpenStudy (anonymous):

If you know Euler's Identities then you can easily verify these: EULER IDENTITIES: \[\large e^{i \theta} = \cos \theta + i.\sin \theta\] \[\large e^{-i \theta} = \cos \theta - i.\sin \theta\] Adding these both the identities: \[\large \cos \theta = \frac{e^{i \theta} + e^{-i \theta}}{2}\] \[\large \sin \theta = \frac{e^{i \theta} - e^{-i \theta}}{2i}\] These are the two identities which will prove your question @UnkleRhaukus , Replace \(\theta\) by \(ix\) you get: \[\large sinix = \frac{e^{i(ix) - e^{-i(ix)}}}{2i} \implies sinix = \frac{e^{-x} - e^x}{2i} \implies sinix = -\frac{e^x - e^{-x}}{2i}\] \[\large \because \frac{-1}{i} = i\] \[\large \therefore \quad sinix = i (\frac{e^x - e^{-x}}{2}) \implies sinix = i.sinhx\] Similarly you can use this for proving \(cosix = coshx\).. Getting @UnkleRhaukus ??

OpenStudy (unklerhaukus):

very nice

OpenStudy (anonymous):

Thanks Uncle Rocks..